# Bridging Textual-Collaborative Gap through Semantic Codes for Sequential Recommendation

### v1

@@ Abstract.
In recent years, substantial research efforts have been devoted to enhancing sequential recommender systems by integrating abundant side information with ID-based collaborative information. This study specifically focuses on leveraging the textual metadata (e.g., titles and brands) associated with items. While existing methods have achieved notable success by combining text and ID representations, they often struggle to strike a balance between textual information embedded in text representations and collaborative information from sequential patterns of user behavior. In light of this, we propose CoCoRec, a novel Code-based textual and Collaborative semantic fusion method for sequential Recommendation. The key idea behind our approach is to bridge the gap between textual and collaborative information using semantic codes. Specifically, we generate fine-grained semantic codes from multi-view text embeddings through vector quantization techniques. Subsequently, we develop a code-guided semantic-fusion module based on the cross-attention mechanism to flexibly extract and integrate relevant information from text representations. In order to further enhance the fusion of textual and collaborative semantics, we introduce an optimization strategy that employs code masking with two specific objectives: masked code modeling and masked sequence alignment. The merit of these objectives lies in leveraging mask prediction tasks and augmented item representations to capture code correlations within individual items and enhance the sequence modeling of the recommendation backbone. Extensive experiments conducted on four public datasets demonstrate the superiority of CoCoRec, showing significant improvements over various sequential recommendation models. Our code is available at https://anonymous.4open.science/r/CoCoRec-6E41.
Sequential Recommendation, Textual-Collaborative Representation Fusion, Semantic Codes
journalyear: 2018copyright: acmcopyrightconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYdoi: 0000000.0000000isbn: 978-1-4503-XXXX-X/18/06price: 15.00ccs: Information systems Recommendation systems
@@ 1. Introduction
Early approaches (Zhang et al., 2019; Hou et al., 2022) achieve moderate success by introducing text representations encoded by pre-trained language models (PLMs) in a straightforward way (e.g., sum and concat), but these methods overlook the semantic gap between ID embeddings and text representations. To address this limitation, some studies adopt various techniques, including contrastive learning (Zhou et al., 2020), attention mechanisms (Liu et al., 2021; Xie et al., 2022b) and graph-based aggregation (Hu et al., 2023), to better fuse ID embeddings and text representations. Furthermore, recent studies propose mapping text representations into semantic codes for sequential recommendation (Petrov and Macdonald, 2024; Hou et al., 2023; Hu et al., 2023) or generative recommendation (Rajput et al., 2023; Zheng et al., 2024). These codes are typically derived through clustering or vector quantization based on text representations (Hua et al., 2023; Rajput et al., 2023). A key merit of this approach is that the shared codes between different items reflect their semantic similarities. Despite their notable effectiveness, these methods struggle to achieve an appropriate trade-off between the textual semantics embedded in text representations and the collaborative semantics implied by sequential patterns. Integrating item ID embeddings with text representations may result in the model overemphasizing textual features while neglecting the collaborative relationships between items (Hou et al., 2023). Additionally, code-based methods rely on code sharing relationships, which alleviates the over-reliance on textual features but inevitably leads to textual semantic loss.
@@
Considering these issues, our idea is to employ semantic codes as a bridge to achieve a better balance between textual information and collaborative information. This approach is motivated by two key characteristics of semantic codes. First, they can be regarded as a specialized form of item IDs, associated with learnable embeddings that can flexibly capture collaborative semantics by modeling interaction sequences. Second, the multiple codes for each item imply semantics of different granularities or levels, injecting prior knowledge of semantic similarity into item representations, which can serve as guidance for refined textual-collaborative semantics integration. To develop our approach, we focus on two aspects: (1) leveraging semantic codes as guidance to flexibly extract useful information from text representations, alleviating the over-reliance on textual information, and (2) learning informative item representations to enhance sequential recommendation.
@@
To this end, we propose CoCoRec, a novel Code-based textual and Collaborative semantic fusion approach for sequential Recommendation. Unlike previous methods that directly utilize text embeddings as item representations, our approach develops a code-guided fusion module to derive semantically enriched item representations. In this module, semantic codes serve as the query to flexibly extract and integrate relevant information from text representations, thereby mitigating the excessive dependence on item text while enabling a more comprehensive learning of code-based collaborative information. To obtain semantic codes with different granularity information, we encode various item attributes (e.g., title, brand, categories) separately as multi-view text representations and quantize them into a tuple of codes through vector quantization. Then, we implement the semantic fusion module as a multi-layer cross-attention network to incorporate code embeddings and multi-view text representations. In order to further enhance semantic fusion and sequence modeling, we propose two optimization objectives through item code masking, namely masked code modeling and masked sequence alignment. For masked code modeling, the semantic fusion module recovers the masked code from multi-view text representations, which has two advantages. Firstly, it can capture the code relationships within individual items. Secondly, it promotes the integration between code embeddings and text representations. For masked sequence alignment, we align the user preference representations derived from interaction sequences involving masked codes with the original one, aiming to enhance the robustness of user preference modeling.
@@
$\bullet$ We propose a novel code-based textual and collaborative semantic fusion method for sequential recommendation, which bridges the textual-collaborative semantic gap through semantic codes.
@@
$\bullet$ Extensive experiments on four public recommendation benchmarks demonstrate the effectiveness of our proposed framework.
Figure 1. Overall framework of CoCoRec, which comprises two key components, i.e., the Semantic-Fusion Module (SFM) and the Sequential Recommendation backbone (SR). SFM stacks Self-Attention (SA) layers, Cross-Fusion (CF) layers, and Feed-Forward Networks (FFN) to enable effective learning of semantic-fused item representations.
@@ 2.1. Problem Statement
In our scenario, we focus on sequential recommendation tasks. Let $\mathcal{U}=\{u_{1},\dots,u_{|\mathcal{U}}|\}$ denote a set of users and $\mathcal{V}=\{v_{1},\dots,v_{|\mathcal{V}|}\}$ denote a set of items. Each item $v$ is associated with a unique item ID and a series of text attributes $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ (e.g., title, brand, category, etc.) where $m$ represents the attribute number. Here, each item attribute $a_{i}$ corresponds to a descriptive text. Additionally, $a_{i}$ is mapped to a tuple of semantic codes, denoted by $C_{i}=(c^{i}_{1},\dots,c^{i}_{k})$, where semantic codes under the same view (e.g., title or brand) share a common code space. Given a user interaction sequence $s=\{v_{1},v_{2},\dots,v_{t}\}$ in chronological order, the objective of sequential recommendation is to predict the next item that the user is most likely to interact with, which is formulated as $P(v_{t+1}|\{v_{1},\dots,v_{t}\})$.
@@ 2.2. Text Embeddings and Semantic Codes
In this part, we describe the process of deriving two distinct types of representations utilized to characterize items. This involves (a) constructing multi-view text embeddings by encoding each item attribute individually, and (b) mapping the text embeddings to semantic codes through multi-level vector quantization methods.
@@ 2.2.1. Multi-View Text Embeddings
Text embeddings are widely acknowledged for encapsulating rich textual information. To effectively harness item content, previous studies (Xu et al., 2024; Liu et al., 2024a; Hu et al., 2023, 2024) typically concatenate all associated text and encode it into a single embedding using pre-trained language models (PLMs). However, it’s difficult to leverage fine-grained item information by jointly encoding all content. Moreover, due to the limited input length of PLMs, the item text often needs to be truncated, leading to potential incomplete semantics. Therefore, to fully exploit attribute-wise semantic information, we propose encoding each item attribute individually to obtain multi-view text embeddings. Formally, given the text attributes $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ of item $v$, the text embedding of attribute $a_{i}$ is formulated as
@@
where $\bm{z}^{t}_{i}$ is the mean pooling representation of the last hidden states of the input.
@@ 2.2.2. Vector Quantization for Semantic Codes
In our framework, semantic codes serve a dual purpose. First, they are linked to learnable embeddings, enabling the flexible capture of collaborative information. Second, they act as bridges to facilitate the fusion of textual and collaborative semantics. Unlike unique IDs, semantic codes allow semantically similar items to share representations, thereby capturing inherent relationships among items. Moreover, the multi-level structure of these codes reflects textual semantics from different aspects, providing guidance for fine-grained semantic fusion. These characteristics make semantic codes particularly effective for item representation, as evidenced by prior research (Hou et al., 2023; Rajput et al., 2023; Petrov and Macdonald, 2024). In our approach, we utilize vector quantization techniques, such as product quantization (PQ) or residual quantization (RQ), to map item attribute embeddings into multi-level discrete codes. Specifically, for an attribute embedding $\bm{z}^{t}_{i}$ of item $v$, the vector quantization process generates a tuple of discrete codes $C_{i}=(c^{i}_{1},\dots,c^{i}_{k})$, where $k$ denotes the number of quantized sub-vectors, which is a pre-defined hyper-parameter. Given that each item has $m$ distinct attributes, this process produces a total of $n_{c}=m\times k$ semantic codes, represented as $(c^{1}_{1},c^{1}_{2},\dots,c^{m}_{k-1},c^{m}_{k})$.
@@ 2.3. Code-Guided Semantic Fusion
To effectively integrate textual and ID-based collaborative semantics, prior research (Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022; Hu et al., 2023; Xu et al., 2024; Liu et al., 2023; Gao et al., 2024) has typically combined text representations with ID embeddings or utilized the discrete codes derived from text embeddings to enhance sequential recommenders. Despite their effectiveness, these methods often struggle to achieve an appropriate trade-off between the textual information encoded in text embeddings and the collaborative information implied by sequential patterns. To address this issue, we propose to employ semantic codes as bridges to extract and integrate relevant information from text representations. A key challenge in this approach lies in fully unleashing the potential of semantic codes to serve as effective bridges. To tackle this challenge, we introduce a code-guided semantic-fusion module that integrates text representations with multi-level semantic code embeddings using cross-attention mechanisms, where semantic codes are considered as the query and text representations are regarded as the key and value. This design enables the semantic codes to flexibly extract and integrate useful information from text representations, mitigating the over-reliance on item text. Subsequently, a Transformer-based recommendation backbone leverages these fused item representations to model sequential user behavior and makes final recommendations.
@@ 2.3.1. Semantic Code Embeddings
Our approach leverages two types of embeddings, i.e., attribute embeddings and code embeddings, to capture textual and collaborative information, respectively. Each item $v$ is associated with $n_{c}$ semantic codes denoted as $(c_{1},\dots,c_{n_{c}})$ (for simplicity, we do not distinguish the attributes). To learn the collaborative correlations among items, we set up $n_{c}$ code embedding tables for lookup. For each position $l$, all items share a common code embedding table $\bm{E}^{(l)}\in\mathbb{R}^{C\times d}$, where $d$ denotes the dimension of the item representations. Using these tables, we look up the code embeddings for item $v$ as $\{\bm{z}^{c}_{1}\}_{i=1}^{n_{c}}$ where $\bm{z}^{c}_{i}\in\bm{E}^{(i)}$. In addition to code embeddings, each item $v$ is also associated with $m$ attribute embeddings, denoted as $\{\bm{z}^{t}_{i}\}_{i=1}^{m}$. Here, $\bm{z}^{t}_{i}\in\mathbb{R}^{d_{e}}$ represents the embedding of the $i$-th attribute, where $d_{e}$ is typically larger than $d$. To ensure compatibility between the two types of embeddings, we apply Principal Component Analysis (PCA) to reduce the dimension of $\bm{z}^{t}_{i}$ to $d$.
@@ 2.3.2. Semantic-Fusion Module
where $\bm{H}^{0}=\bm{Z}^{c}$ and $L$ is the number of layers in the semantic-fusion module and $\operatorname{FFN}(\cdot)$ denotes the multilayer perceptron network. The ordering of elements within these two sequences holds no intrinsic meaning. Therefore, we omit the inclusion of positional encoding to neglect sequential relationships. The final semantically fused sequence representation, $\bm{H}\in\mathbb{R}^{n_{c}\times d}$, is aggregated to generate the semantic-fused item representation $\bm{e}=\operatorname{Pool}\left(\bm{H}\right)$, where $\operatorname{Pool}(\cdot)$ denotes mean pooling. To enhance semantic code embedding learning, the pooled semantic code representation is integrated, yielding the final item representation $\tilde{\bm{e}}=\bm{e}+\operatorname{Pool}\left(\bm{Z}^{c}\right)$.
@@ 2.4. Enhanced Representation Learning via Code Masking
Compared with existing semantic fusion methods (Zhang et al., 2019; Hu et al., 2023; Xu et al., 2024), optimizing our framework is more challenging due to its two hierarchically arranged components. A single recommendation optimization objective is insufficient to ensure the comprehensive learning of the entire framework, particularly for the code representations. To address this issue, we propose two optimization objectives based on code masking to further enhance the representation learning in CoCoRec. Specifically, we introduce the Masked Code Modeling (MCM) task, which promotes the semantic fusion between text and codes to refine item representations. Additionally, we incorporate the Masked Sequence Alignment (MSA) task to improve the sequence modeling capabilities of the recommendation backbone.
@@
Masked Sequence Alignment. MCM is primarily used to optimize the SFM. In order to enhance the sequence representation learning of the recommendation backbone, we introduce the Masked Sequence Alignment (MSA) task, using the contrastive learning technique that have been widely validated in sequential recommendation tasks (Xie et al., 2022a; Chen et al., 2022b; Qiu et al., 2022; Ye et al., 2023). In MSA, code masking serves as an effective data augmentation strategy to generate augmented item representations. This approach increases the diversity of inputs to the recommendation backbone, thereby refining its ability to model user sequences. Specifically, we first generate augmented item representations by feeding the masked code sequence and associated text representations into the SFM. These augmented item representations are then passed to the recommendation backbone to produce augmented sequence representations. Subsequently, the two sequence representations are aligned using the InfoNCE loss. The representations derived from the same sequence are considered as positive samples, while those from different sequences within the same batch are treated as negative samples. The objective function for masked sequence alignment is formulated as follows:
@@
where $\mathcal{B}_{\bm{r}}$ and $\mathcal{B}_{\dot{\bm{r}}}$ denote the batch of original sequence representations and augmented sequence representations generated by the recommendation backbone, respectively. $\operatorname{InfoNCE}(\cdot,\cdot,\cdot)$ denotes the InfoNCE loss, which can be formulated as:

### v2

@@ Abstract.
In recent years, substantial research efforts have been devoted to enhancing sequential recommender systems by integrating abundant side information with ID-based collaborative information. This study specifically focuses on leveraging the textual metadata (e.g., titles and brands) associated with items. While existing methods have achieved notable success by combining text and ID representations, they often struggle to strike a balance between textual information embedded in text representations and collaborative information from sequential patterns of user behavior. In light of this, we propose CCFRec, a novel Code-based textual and Collaborative semantic Fusion method for sequential Recommendation. The key idea behind our approach is to bridge the gap between textual and collaborative information using semantic codes. Specifically, we generate fine-grained semantic codes from multi-view text embeddings through vector quantization techniques. Subsequently, we develop a code-guided semantic-fusion module based on the cross-attention mechanism to flexibly extract and integrate relevant information from text representations. In order to further enhance the fusion of textual and collaborative semantics, we introduce an optimization strategy that employs code masking with two specific objectives: masked code modeling and masked sequence alignment. The merit of these objectives lies in leveraging mask prediction tasks and augmented item representations to capture code correlations within individual items and enhance the sequence modeling of the recommendation backbone. Extensive experiments conducted on four public datasets demonstrate the superiority of CCFRec, showing significant improvements over various sequential recommendation models. Our code is available at https://github.com/RUCAIBox/CCFRec.
Sequential Recommendation; Textual-Collaborative Representation Fusion; Semantic Codes
✉ Corresponding author.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736869isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Recommendation systems
@@ 1. Introduction
Early approaches (Zhang et al., 2019; Hou et al., 2022) achieve moderate success by introducing text representations encoded by pre-trained language models (PLMs) in a straightforward way (e.g., sum and concat), but these methods overlook the semantic gap between ID embeddings and text representations. To address this limitation, some studies adopt various techniques, including contrastive learning (Zhou et al., 2020), attention mechanisms (Liu et al., 2021; Xie et al., 2022b) and graph-based aggregation (Hu et al., 2023), to better fuse ID embeddings and text representations. Furthermore, recent studies propose mapping text representations into semantic codes for sequential recommendation (Petrov and Macdonald, 2024; Hou et al., 2023; Hu et al., 2023) or generative recommendation (Rajput et al., 2023; Zheng et al., 2024; Hou et al., 2025). These codes are typically derived through clustering or vector quantization based on text representations (Hua et al., 2023; Rajput et al., 2023). A key merit of this approach is that the shared codes between different items reflect their semantic similarities. Despite their notable effectiveness, these methods struggle to achieve an appropriate trade-off between the textual semantics embedded in text representations and the collaborative semantics implied by sequential patterns. Integrating item ID embeddings with text representations may result in the model overemphasizing textual features while neglecting the collaborative relationships between items (Hou et al., 2023). Although code-based methods can alleviate the over-reliance on text, they suffer from textual semantic loss in a certain extent.
@@
Considering these issues, our idea is to employ semantic codes as a bridge to achieve a better balance between textual and collaborative information. This approach is motivated by two key characteristics of semantic codes. First, they can be regarded as a specialized form of item IDs, associated with learnable embeddings that can flexibly capture collaborative semantics by modeling interaction sequences. Second, the multiple codes for each item imply semantics of different granularities or levels, injecting prior knowledge of semantic similarity into item representations, which can serve as guidance for refined textual-collaborative semantics integration. To develop our approach, we focus on two aspects: (1) leveraging semantic codes as guidance to flexibly extract useful information from text representations, alleviating the over-reliance on textual information, and (2) learning informative item representations to enhance sequential recommendation.
@@
To this end, we propose CCFRec, an innovative Code-based textual and Collaborative semantic Fusion approach for sequential Recommendation. Unlike previous methods that directly utilize text embeddings as item representations, our approach develops a code-guided fusion module to derive semantically enriched item representations. In this module, semantic codes serve as queries to flexibly extract relevant information from text representations, thereby mitigating the excessive dependence on item text while enabling a more comprehensive learning of code-based collaborative information. To obtain semantic codes with different granularity information, we encode various item attributes (e.g., title, brand, categories) separately as multi-view text representations and quantize them into a tuple of codes through vector quantization. Then, we implement the semantic fusion module as a multi-layer cross-attention network to incorporate code embeddings and multi-view text representations. In order to further enhance semantic fusion and sequence modeling, we propose two optimization objectives through item code masking, namely masked code modeling and masked sequence alignment. For masked code modeling, the semantic fusion module recovers the masked code from multi-view text representations, which has two advantages. Firstly, it can capture the code relationships within individual items. Secondly, it promotes the integration between code embeddings and text representations. For masked sequence alignment, we align the user preference representations derived from interaction sequences involving masked codes with the original one, aiming to enhance the robustness of user preference modeling.
@@
$\bullet$ We propose CCFRec, a novel code-based textual and collaborative semantic fusion method for sequential recommendation, which bridges the textual-collaborative gap through semantic codes.
@@
$\bullet$ Extensive experiments on four public recommendation benchmarks demonstrate the effectiveness of our proposed framework CCFRec.
Figure 1. Overall framework of CCFRec, which comprises two key components, i.e., the Semantic-Fusion Module (SFM) and the Sequential Recommendation backbone (SR). SFM stacks Self-Attention (SA) layers, Cross-Fusion (CF) layers, and Feed-Forward Networks (FFN) to enable effective learning of semantic-fused item representations.
@@ 2.1. Problem Statement
In our scenario, we focus on sequential recommendation tasks. Let $\mathcal{U}=\{u_{1},\dots,u_{|\mathcal{U}}|\}$ denote a set of users and $\mathcal{V}=\{v_{1},\dots,v_{|\mathcal{V}|}\}$ denote a set of items. Each item $v$ is associated with a unique item ID and a series of text attributes $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ (e.g., title, brand, category, etc.) where $m$ represents the attribute number. Here, each item attribute $a_{i}$ corresponds to a descriptive text. Additionally, $a_{i}$ is mapped to a tuple of semantic codes, denoted by $C_{i}=(c^{i}_{1},\dots,c^{i}_{k})$, where semantic codes under the same view (e.g., title or brand) share a common code space. Given a user interaction sequence $s=\{v_{1},v_{2},\dots,v_{t}\}$ in chronological order, the objective of sequential recommendation is to predict the next item that the user is most likely to interact with, which is formulated as $\max P(v_{t+1}|\{v_{1},\dots,v_{t}\})$.
@@ 2.2. Text Embeddings and Semantic Codes
In this part, we describe the process of deriving two distinct representations for items. This involves (a) constructing multi-view text embeddings by encoding each item attribute individually, and (b) mapping the text embeddings to semantic codes through multi-level vector quantization methods.
@@ 2.2.1. Multi-View Text Embeddings
Text embeddings are widely acknowledged for encapsulating rich textual information. To effectively harness item content, previous studies (Xu et al., 2024; Liu et al., 2024a; Hu et al., 2023, 2024) typically concatenate all associated text and encode it into a single embedding using pre-trained language models (PLMs). However, it’s difficult to leverage fine-grained item information by jointly encoding all content. Moreover, due to the limited input length of PLMs, the item text often needs to be truncated, leading to potential incomplete semantics. Therefore, to fully exploit attribute-wise semantic information, we propose encoding each item attribute individually to obtain multi-view text embeddings. Formally, given text attributes $\mathcal{A}=\{a_{1},\dots,a_{m}\}$ of item $v$, the text embedding of attribute $a_{i}$ is formulated as
@@
where $\bm{z}^{t}_{i}$ denotes the mean pooling representation of the output.
@@ 2.2.2. Vector Quantization for Semantic Codes
In our framework, semantic codes serve a dual purpose. First, they are linked to learnable embeddings, enabling the flexible capture of collaborative information. Second, they act as bridges to facilitate the fusion of textual and collaborative semantics. Unlike unique IDs, semantic codes allow semantically similar items to share representations, thereby capturing inherent relationships among items. Moreover, the multi-level structure of these codes reflects textual semantics from different aspects, providing guidance for fine-grained semantic fusion. These characteristics make semantic codes particularly effective for item representation, as evidenced by prior research (Hou et al., 2023; Rajput et al., 2023; Petrov and Macdonald, 2024). In our approach, we utilize vector quantization techniques, such as product quantization (PQ) and residual quantization (RQ), to map item attribute embeddings into multi-level discrete codes. Specifically, for an attribute embedding $\bm{z}^{t}_{i}$ of item $v$, the vector quantization process generates a tuple of discrete codes $C_{i}=(c^{i}_{1},\dots,c^{i}_{k})$, where $k$ denotes the number of quantized sub-vectors, which is a pre-defined hyper-parameter. Given that each item has $m$ distinct attributes, this process produces a total of $n_{c}=m\times k$ semantic codes, represented as $(c^{1}_{1},c^{1}_{2},\dots,c^{m}_{k-1},c^{m}_{k})$.
@@ 2.3. Code-Guided Semantic Fusion
To effectively integrate textual and ID-based collaborative semantics, prior research (Zhang et al., 2019; Zhou et al., 2020; Hou et al., 2022; Hu et al., 2023; Xu et al., 2024; Liu et al., 2023; Gao et al., 2024) has typically combined text representations with ID embeddings or utilized the discrete codes derived from text embeddings to enhance sequential recommenders. Despite their effectiveness, these methods often struggle to achieve an appropriate trade-off between the textual information encoded in text embeddings and the collaborative information implied by sequential patterns. To address this issue, we propose to employ semantic codes as bridges to extract and integrate relevant information from text representations. A key challenge in this approach lies in fully unleashing the potential of semantic codes to serve as effective bridges. To tackle it, we introduce a code-guided semantic-fusion module that integrates text representations with multi-level semantic code embeddings using cross-attention mechanisms, where semantic codes are considered as the query and text representations are regarded as the key and value. This design enables the semantic codes to flexibly extract useful information from text representations, mitigating the over-reliance on item text. Subsequently, a Transformer-based recommendation backbone leverages these fused item representations to model sequential user behavior and makes final recommendations.
@@ 2.3.1. Semantic Code Embeddings
Our approach leverages two types of embeddings, i.e., attribute embeddings and code embeddings, to capture textual and collaborative information, respectively. Each item $v$ is associated with $n_{c}$ semantic codes denoted as $(c_{1},\dots,c_{n_{c}})$ (for simplicity, we do not distinguish the attributes). To learn the collaborative correlations among items, we set up $n_{c}$ code embedding tables for lookup. For each position $l$, all items share a common code embedding table $\bm{E}^{(l)}\in\mathbb{R}^{C\times d}$, where $d$ denotes the dimension of the item representations. Using these tables, we look up the code embeddings for item $v$ as $\{\bm{z}^{c}_{i}\}_{i=1}^{n_{c}}$ where $\bm{z}^{c}_{i}\in\bm{E}^{(i)}$. In addition to code embeddings, each item $v$ is also associated with $m$ attribute embeddings, denoted as $\{\bm{z}^{t}_{i}\}_{i=1}^{m}$. Here, $\bm{z}^{t}_{i}\in\mathbb{R}^{d_{e}}$ represents the embedding of the $i$-th attribute, where $d_{e}$ is typically larger than $d$. To ensure compatibility between two types of embeddings, we apply Principal Component Analysis (PCA) to reduce the dimension of $\bm{z}^{t}_{i}$ to $d$.
@@ 2.3.2. Semantic-Fusion Module
where $\bm{H}^{0}=\bm{Z}^{c}$ and $L$ is the number of layers in the semantic-fusion module and $\operatorname{FFN}(\cdot)$ denotes the multilayer perceptron network. The ordering of elements within these two sequences holds no intrinsic meaning. Therefore, we omit position embeddings to neglect sequential relationships. The final semantically fused sequence representation, $\bm{H}\in\mathbb{R}^{n_{c}\times d}$, is aggregated to generate the semantic-fused item representation denoted as $\bm{e}=\operatorname{Pool}\left(\bm{H}\right)$, where $\operatorname{Pool}(\cdot)$ denotes mean pooling. To enhance semantic code embedding learning, the pooled semantic code representation is integrated, yielding the final item representation $\tilde{\bm{e}}=\bm{e}+\operatorname{Pool}\left(\bm{Z}^{c}\right)$.
@@ 2.4. Enhanced Representation Learning via Code Masking
Compared with existing semantic fusion methods (Zhang et al., 2019; Hu et al., 2023; Xu et al., 2024), optimizing our framework is more challenging due to its two hierarchically arranged components. A single recommendation optimization objective is insufficient to ensure the comprehensive learning of the entire framework, particularly for the code representations. To address this issue, we propose two optimization objectives based on code masking to further enhance the representation learning in CCFRec. Specifically, we introduce the Masked Code Modeling (MCM) task, which promotes the semantic fusion between text and codes to refine item representations. Additionally, we incorporate the Masked Sequence Alignment (MSA) task to improve the sequence modeling capabilities of the recommendation backbone.
@@
Masked Sequence Alignment. MCM is primarily used to optimize the SFM. In order to enhance the sequence representation learning of the recommendation backbone, we introduce the Masked Sequence Alignment (MSA) task, using the contrastive learning technique that has been widely validated in sequential recommendation tasks (Xie et al., 2022a; Chen et al., 2022b; Qiu et al., 2022; Ye et al., 2023). In MSA, code masking serves as an effective data augmentation strategy to generate augmented item representations. This approach increases the diversity of inputs to the recommendation backbone, thereby refining its ability to model user sequences. Specifically, we first generate augmented item representations by feeding the masked code sequence and associated text representations into the SFM. These augmented item representations are then passed to the recommendation backbone to produce augmented sequence representations. Subsequently, the two sequence representations are aligned using the InfoNCE loss. The representations derived from the same sequence are considered as positive samples, while those from different sequences within the same batch are treated as negative samples. The objective function for masked sequence alignment is formulated as follows:
@@
where $\mathcal{B}_{\bm{r}}$ and $\mathcal{B}_{\dot{\bm{r}}}$ denote the batch of original sequence representations and augmented sequence representations generated by the recommendation backbone, respectively. $\operatorname{InfoNCE}(\cdot,\cdot,\cdot)$ denotes the InfoNCE loss, which can be written as: